The length of a rectangle is 9 centimeters more than four times the width of the area of the rectangle is 198 centimeters squared find the length and the width
Answer by tiffany222(56)

Let width be x
Let length be y
y = 4x + 9
A = l * w = x * y
A = x (4x + 9)
198 = 4x^2 + 9x
0 = 4x^2 + 9x - 198
(x - 6) (4x + 33) = 0
x = 6, -33/4
x = 6 only because -33/4 is rejected
Substituting x to find y:
y = 4x + 9
y = 4 (6) + 9
y = 24 + 9
y = 33
Final Answer:
Width = 6 cm
Length = 33 cm